### v1

@@ 4.2 RQ2: User Discussions in App Reviews
4.2.1 Approach
To understand the topics discussed by users of disaster apps, we employed BERTopic [36], a topic modeling technique, on our dataset of 28,161 reviews. We then performed manual cluster merging and naming, followed by cluster refinement to remove outliers. The process of topic modeling can be seen in Figure 7, and the details of the approach are explained below.
Figure 7: Topic Modelling Process
BERTopic: In this research, BERTopic was utilized for topic modeling. BERTopic converts documents into vectors using the Sentence-BERT (SBERT) framework. SBERT is utilized for document embedding because it can retain semantic information. BERTopic provides two options for sentence embedding: all-MiniLM-L6-v2 for English sentences and paraphrase-multilingual-MiniLM-L12-v2 for multilingual sentences [37]. The first option was chosen because we translated all non-English reviews into English. Subsequently, the dimensionality of the embedding vector is reduced using UMAP (Uniform Manifold Approximation and Projection) [38].
UMAP was selected for dimension reduction in BERTopic for its superior embedding stability and significantly faster runtime compared to other methods such as T-SNE [39]. Once reduced, the embedding vector is clustered using HDBSCAN (Hierarchical density-based clustering) [40]. HDBSCAN was chosen because this model can prevent unrelated documents (outliers) from being assigned to any cluster and is more robust in parameter selection. HDBSCAN has parameters that need to be specified, among which is min_topic_size [37]. This parameter ensures that each cluster has at least min_topic_size members. If there are numerous points in close proximity but fewer than the min_topic_size, these points will either join another cluster or become outliers. In this study, the min_topic_size used in this clustering was 100.
Cluster Merging and Naming: After the model training is complete, BERTopic will provide results in the form of topic IDs, topic names, keywords, and the number of members in each topic. To ensure the accuracy of the topic/cluster naming, a manual analysis is conducted to rename each topic. This analysis involves examining the keywords of each topic, as well as checking 50 data points that are closest in distance to the center of each cluster. After all clusters are appropriately named, similar clusters are merged, for example, clusters named “app notification” and “app alert” are considered the same topic but differ only in terminology.
Cluster Refinement: In the topic modeling process, each identified topic is refined by removing reviews that don’t align well with the topic’s main theme. This involves using BERTopic to re-cluster the reviews within a topic into two groups: outliers and non-outliers. Reviews classified as outliers are discarded, ensuring that the remaining, relevant reviews are retained and become the final result of the topic modeling process. If a topic contains no outliers, the quantity of reviews remains unchanged.
4.2.2 Results
Our topic modeling process using BERTopic resulted in 20 topics. These topics are sorted based on the number of reviews per topic in Table 6. Each topic has a minimum of 50 reviews, as specified by the min_topic_size parameter in the BERTopic model. In the following section, we will briefly discuss each topic along with examples.
@@ 1 Alert Functionality alert, notification, warning, report 3,129 17
2 App Satisfaction good, great, nice, love, like 2,194 10
@@ 4.2.3 Identified Topics
However, there are some reviews in which users were disappointed with the alert functionality. One user complained about the late delivery of alerts “… I didn’t receive timely warnings. Alerts can be sent between 10 minutes to an hour late after an earthquake occurs!”
@@
2) App Satisfaction: The second most discussed topic is app satisfaction, with 2,194 reviews, where users indicate they are generally happy with the app (see Table 6). The majority of reviews in this topic are brief expressions of satisfaction. The median word count for reviews in this topic is only 10, the lowest among all topics. An example of a short review is “Great app, easy to use, and provides good information.” However, there are also longer reviews where users explain their experience using the app, such as “I really appreciated how up-to-date the app is with fire information! The other day when I was driving on back roads in country Victoria, I saw smoke and checked the app, and realized I was heading straight for a quickly growing new grass fire.”
@@
3) Map Usage: The third most-discussed topic is regarding map usage, with 567 reviews (see Table 6. Although only 21 out of 45 disaster apps had a map feature (see Table 1), this topic has many reviews, many of them negative. The common complaints are level of details on the map, clumping of alerts when zooming and issues in map functionality and usability, such as “… The UI is incredibly clunky, maps are extremely slow to load, and warning icons disappear when moving across the map…”
@@
However, some users were satisfied with the map feature in the app. They are happy about aspects such as accuracy of the maps, fastness in loading, ease of reading and usage. This is shown in the review “… The map feature works very well, and zooming in provides great detail. Positioning the map for a particular global view also works very well…”
@@
4) Monetary Transactions and Ads: This topic contains customer reviews regarding subscriptions, payments, and advertising. Some users give positive ratings and comments because the app is free and does not have ads, such as “Brilliant. This is actual free app… It’s the first 5-star app for me… Plus, there are no ads!”
@@
However, many other users complain about the disruptive ads, for example “… with the latest update, you don’t just get banner ads,but you also get forced full-page pop-up ads, which are incredibly obtrusive and annoying.” Other users also complained about dissatisfaction after purchasing the app or upgrading to the pro version, such as “I paid for the Pro version, and there’s nothing special about it. The app doesn’t warn you right away; it has like a 10-15 minute wait time before it pops up. Living in Alaska, this app is worthless to me.”
@@
5) Location Functionality: This topic contains user reviews related to location functionality, such as satisfaction regarding the ability to add their preferred location to be monitored. This is shown in review “I have friends and family scattered all over the country. The app allows me to add the places where they live and provides alerts if there are hazards such as wildfire weather, earthquakes, or hurricanes approaching. It’s easy to customize the alerts I want and to add new places. It’s also great that it will monitor wherever I am.”
@@
Several other reviews mention user dissatisfaction due to inability to add locations, location errors, or inability to find locations. One such review is “The app will not allow me to add a home location or turn on live monitoring, rendering the entire app useless. My location permissions are set to ’always’, and the app can detect my location via GPS. However, saving never works and instead always produces an error message.”
@@
6) Disaster Updates: This topic contains user reviews from individuals who are pleased to stay informed about disasters occurring in various parts of the world. This topic only contained short reviews. The median word count for user reviews in this topic is the second lowest, with only 17 words (See Table 6). Some examples are “Great to be able to look all over the world and see what is happening!”, “…The information it contains offers a full view of what earth is doing. Checking out this app is like watching the earth breath and grow or gag and stumble…”. and “… Since I have friends all around the world, I like knowing where the earth is quaking. At times, it is shocking to know how many tremors and earthquakes happen everyday, but I would rather know, than not.”
@@
7) App Crashes: This topic contains user reviews related to the app being unusable due to crashes. Almost all reviews are negative. App crashes can occur in different stages of using the app. One instance is when users attempt to open the application “… lately, either it won’t open at all, or it closes after a couple of seconds…”. It can also happen in other stages of using the app such as “… Sometimes it crashes when I am in the settings, and I have to restart it…”
@@
8) Alert Sound. This topic discusses issues in the level of alert sounds. Sometimes the issue is with the loudness of the alert sound “… but the notification sound is loud and annoying. I turned down my notification volume on my phone, which made other apps quieter, not what I wanted. However, within 2 days, the volume for this app was so high-pitched that it hurts your ears, and the volume on other apps was still low.”. Some users had complaints about the absence of sound, causing users to be unaware if a disaster occurs “A tornado alert was issued by this app, but there are no sounds at all! A wind chime would be more useful than this app. An audible alarm that alerts you should be mandatory on all weather apps.”
9) App Dissatisfaction. This topic discusses user disappointment with the application, with all reviews negative. Disappointment typically contains issues such as lack of developer feedback, issues in functionalities like logging in. The following review shows a combination of these issues plus a complaint about not receiving benefits after upgrading to the pro version. “I had trouble with the free version not staying logged in, so I bought the subscription hoping it would stay connected so I could receive emergency alerts. But it made no difference; the app still won’t stay logged in. It’s useless as an Emergency Notification app. I tried emailing for assistance twice, but each time, I received a reply stating that the mailbox was too full for delivery. It must be full of other frustrated users who are getting no help from the developers. VERY DISAPPOINTED!”
@@
10) Slow Performance: This topic contains user reviews related to slow app performance experiences. It can involve slow loading features or slowness when using the features. Example reviews include “Worst app. If you want to view the map and look at details, you have to wait 10 minutes. Super slow.” and “… it’s very slow trying to move around the app, especially when trying to make a watch list…”
@@
11) Bug Fixes: This topic contains customer reviews regarding bug fixing, including both request to fix bugs and appreciation about bugs that have been resolved. Example of a request to fix bugs is “I have not received any notifications from this app for months! I have logged out, uninstalled, and reinstalled multiple times, but still nothing… Please fix it!”. Example of an appreciation about a bug fix is “Thank you very much for fixing the app! It’s the best one for looking up earthquakes and has a great interface and filters!”
@@
12) App Configuration: This topic contains user reviews related to app configuration. Users may express satisfaction with the flexible configuration, or they may encounter problems with the application configuration. Satisfied customer review can look like this “… a fantastic tool. It offers many settings for configurations tailored to specific needs.” Problems in configuration can be “It won’t open settings when I click, and I can’t adjust which warnings or emergencies I want alerts for.”
@@
13) User Interface: This topic contains user reviews related to the app’s interface. Many users may express satisfaction with the UI such as “Great clean new look. Looking forward to future updates, more new features, and improved design.”. They also highlight problematic issues in UI aspects, such as color and size. One review particularly from a user with vision issues stated “The color combination of black with dark blue is terrible. I would have given it five stars if I could actually read some of the menu labels at the bottom of the screen. My eyesight is limited to one eye, with about 65 to 70% vision out of that eye. I need good contrast to be able to see clearly.”
@@
14) User Safety: This topic contains reviews from users who are happy because they feel safe and helped by this application. This topic often contains lengthy reviews as users share their experiences using this app. The median word count per user review in this topic is 22, making it one of the third highest among other topics (See Table 6). An example reviews are “… puts a lot of safety ideas for different emergencies in one place. I really like the monitoring feature, with family scattered or traveling a lot” and “Last night we had an earthquake 40km away had this app on my phone unopened for many months but still got a notification seconds before we felt it. It allowed us to immediately recognise what was happening and start heading to a safe area. Quite impressed!”
@@
15) Sign Up and Sign In Problems: This topic contains user reviews complaining about not being able to register or problems during the login process, with almost all reviews negative. Examples are “Keeps saying registration failed… Already missed a fire near me due to this new app. Please fix.” and “It won’t even let me log into the app at all. It keeps saying my login credentials are incorrect, but when I use the same login info on the web browser, it logs right in. I uninstalled it and won’t be reinstalling it!!”
@@
16) Network Connectivity: This topic contains user complaints regarding the inability to use the app due to connectivity issues, almost all reviews negative. For example, as shown in “Keeps telling me no internet connection detected. Yet I can do everything else and use all other apps without any problems.” Another example of a problem is a user failing to connect to the app server “Many times I get an error message about not being able to connect to the server…”. Some users requested the offline mode for the app as shown in review “Great app, although it’s pretty limited without power or internet connectivity. It would be better if it could be mirrored locally for offline use.”
@@
17) Watch Zone Functionality: This topic contains the longest reviews compared to other topics, with the median word count being 27 (see Table 6). This topic contains complaints about the ‘watch zone’ not being able to be set such as “Almost useless. You cannot create watch zones. It just says ’unable to create watch zones’ over and over…”. Besides that, this topic also contains complaints about the watch zone not working properly as shown in review “It gives me notifications every time there’s a thunderstorm anywhere in the state, but when there was actually a fire in my watch zone, it didn’t come up.”
@@
18) Battery Drainage: This topic is quite straightforward, covering the issue of the app consuming a lot of battery, with almost all reviews highly negative. “… the app was using a huge amount of battery and data in the background. Thirty-one percent battery drain was attributed to this app, along with 337MB of data in only a few days of background activity.”
@@
19) Reliability: This topic contains customer reviews regarding whether users find the app reliable or not. Example positive perceptions look like e.g. “The app is very reliable and trustworthy. I depend on it as I live in a quake-prone area, and the information that this app provides is very important to me and my family.” A negative perception is reflected in the review “Unreliable for personal safety and potentially life-threatening during the recent 2020 fires for many people. I don’t know if the networks were jammed or if the app was broken, but we had to rely on looking out a window for a true assessment of whether there are fires near me or not.”
@@
20) Language Preferences: This topic contains customer reviews related to language issues. Many users wish for the app to be available in a language they understand or are fluent, as reflected in the review “… The only criticism I see right now is the lack of supported languages. I think an English version should at least be provided. Other languages could be added by volunteers, for example, through open-source contributions or via a "donate a translation" button in the app.” Some users suggested alternatives for in app translation : “Could you make it possible to have the English translation? Or at least make it possible to copy the text so that I can translate it.”
4.2.4 Summary
Comparison of Disaster App Topics with other Apps: From the topics generated by our model, it is evident that there are two special features exclusive to disaster-related applications which are not found in other types of apps. These include Alert and Watch zone functionalities.
1) Alert Feature: This is considered one of the most important features in disaster-related applications because they enable users to take immediate action to ensure their safety before or when a disaster strikes.
2) The Watch Zone Feature: This allows users to set a specific radius for receiving alerts, ensuring they are not overwhelmed by unnecessary notifications and are only informed about disasters near their location. Additionally, the watch zone can be customized to monitor specific locations, not just those within a certain radius. This feature is particularly useful for individuals with family or friends living in particular areas and wishing to ensure their safety.
Distribution of Topics Across Reviews: An analysis of Table 6 reveals that the topics of alert functionality and app satisfaction, with 3,129 and 2,194 reviews, significantly outnumber the mentions of other topics. These two topics combined constitute 48.6% of the total reviews, nearly half of the contributions across all topics. This substantial percentage highlights the prominence of app satisfaction, suggesting that at least 20% of the users express satisfaction and commend the app’s performance. In contrast, the topic of alert functionality presents a more complex picture. It is challenging to draw definitive conclusions about user satisfaction from this topic alone. Complaints may arise from various issues, such as delayed alerts, failure to receive alerts, excessive alerts, or reliability concerns. Conversely, satisfaction could stem from timely alerts that effectively precede disasters or alerts that deliver crucial information. Hence, we defined a RQ3 to better understand the users’ positive or negative sentiments towards these topics.
@@ 4.3 RQ3: Challenges in Disaster Apps
4.3.1 Approach
Figure 8: Sentiment Analysis Process
Sentiment Analysis: This RQ examines which topics have the highest percentage of negative reviews. This information can help developers identify dominant user complaints and prioritize their resolution. Ratings serve as an indicator to assess user satisfaction with the app. However, there is potential for Text-Rating-Inconsistency (TRI) to arise because some users may give ratings carelessly [41]. For example, they may give a positive rating immediately after downloading the app without trying it. Another example is a user who provides numerous complaints but assigns a middle rating (a rating of 3). Therefore, relying solely on one indicator may not be robust to ensure that the review aligns with the rating.
We use VADER (A Parsimonious Rule-based Model for Sentiment Analysis of Social Media Text) [42] to predict sentiment in user reviews. The process of sentiment analysis is shown in Figure 8. The outcome of VADER includes four types of sentiment scores: negative scores, neutral scores, positive scores, and compound scores. VADER considers that each sentence may contain both negative and positive elements. For example, consider the review “Currently, I like the app very much, but you should be able to change the language within the app.” This review contains a positive element: “I like the app very much” but also a negative element: “you should be able to change the language”.
VADER provides a combined score of these elements, known as a compound score, in addition to the individual sentiment scores. This compound score, ranging from -1 to 1, will be used as the result of sentiment analysis in this study. We will classify sentiment as positive (score >= 0.05), neutral (0.05 > score > -0.05), or negative (score <= -0.05), following similar approaches used in prior work [31].
Percentage of negative reviews for each topic: In this analysis, we examine the percentage of negative reviews for each topic based on sentiment from VADER or rating (with a value of less than 2). Next, we identify which topics have the highest percentage of negative reviews in terms of rating and sentiment analysis, enabling developers to prioritize these topics for improvement.
4.3.2 Results
@@
Negative   3834 [13.7%]
Neutral   1911 [4.4%]
Positive   6383 [21%]
@@
Negative   3834 [9.5%]
Neutral   1233 [6.8%]
Positive   5890 [22.8%]
Distribution of review sentiment: Table 7 shows the distribution of sentiment analysis across our user review dataset. The percentage of positive sentiment from the user ratings is 21%, and from the sentiment score is 22.8%. The neutral percentage in the user rating is 4.4%, slightly lower than the sentiment score, which is 6.8%. The negative percentage of users is 13.7%, while for the sentiment score, it is 9.5%. From this, it can be seen that the model assumes there are some negative user reviews, but they are actually in the positive and neutral categories, accounting for around 1% and 3%, respectively. However, in general, the distribution of sentiment is the same.
Table 8: The percentage of negative reviews and the average score is based on sentiment analysis and user ratings.
Topic Percentage
Negative
Sentiment
Review
Average
Converted
Sentiment
Score
Percentage
Negative
Rating
Score
Average
Rating
Score
App Dissatisfaction 56.67% 2.70 77.46% 1.85
Sign Up and Sign In Problems 40.00% 2.94 80.65% 1.65
Network Connectivity 37.14% 2.99 62.86% 2.28
App Configuration 32.11% 3.18 42.28% 2.89
Watch Zone Functionality 31.65% 3.23 53.24% 2.53
Slow Performance 31.13% 3.27 44.90% 2.86
User Safety 30.37% 3.55 30.84% 3.71
Alert Sound 30.09% 3.33 29.00% 3.47
Monetary Transactions and Ads 29.52% 3.40 45.39% 3.01
Battery Drainage 28.68% 3.25 57.35% 2.47
Location Functionality 26.12% 3.34 42.35% 2.86
Language Preferences 25.71% 3.37 44.29% 2.89
Reliability 25.71% 3.33 44.76% 3.17
Bug Fixes 24.91% 3.59 45.61% 2.85
App Crashes 24.44% 3.24 68.02% 2.12
Alert Functionality 21.96% 3.47 30.17% 3.48
App Satisfaction 18.87% 3.60 21.01% 3.97
Map Usage 18.52% 3.64 29.63% 3.48
User Interface 17.62% 3.75 17.18% 3.81
Keep Updates 5.81% 3.86 4.21% 4.66
Highest Percentage of Negative Reviews Topics:Table 8 shows the percentage of negative reviews generated from sentiment analysis. Not surprisingly, the topic with the highest percentage is app dissatisfaction. Additionally, it ranks second highest based on user ratings, which is expected given the topic’s nature involving user complaints, either regarding specific features or more general aspects. The next topics with the highest negative reviews based on sentiment analysis are Sign Up and Sign In Problems at 40%, Network Connectivity at 37.14%, and App Configuration at 32.11%. Meanwhile, for user ratings, the highest percentage of negative reviews were found in the topics of Sign Up and Sign In Problems at 80.65%, App Crashes at 68.02%, and Network Connectivity at 62.86%. In terms of average converted sentiment score and average rating, these topics generally have a value below 3, indicating dissatisfaction among users.
5 Recommendations

### v2

@@ 4.2 RQ2: User Discussions in App Reviews
To determine the optimal number of clusters, we evaluated using the Dunn Index and Davies-Bouldin Index. A higher Dunn Index indicates that the clusters are more well-separated, whereas a lower Davies-Bouldin Index indicates better clustering results. We tested a range of topics from 10 to 30. After the experiments, the results are shown in Figure 8. From this figure, it can be seen that the optimal number of clusters is 22.
Figure 8: Evaluation Metrics of BERTopic
However, the automatic topic names generated by BERTopic were not always representative, such as ‘interface_features_ui_layout’ and ‘fix_please_issues_problems’. We therefore renamed these clusters as User Interface and Bug Fixes. After all clusters were appropriately named, similar clusters were merged; for example, clusters named “app notification” and “app alert” were considered the same topic.
Following this merging process, we obtained 18 distinct topics. These topics were then ranked according to the number of reviews per topic, as shown in Table 6. In that table, we provide the topic ID, topic name, keywords representing each topic, total number of reviews in the topic, and review length. The review length in the table refers to the median number of words per user review. We use the median instead of the mean because it is more robust for skewed distributions.
Among the identified topics listed in Table 6, two features—the Alert feature and the Watch Zone feature—stand out as unique to disaster-related applications and are not observed in other types of apps. Notably, the alert functionality is the most frequently discussed theme. A detailed discussion of each topic generated by BERTopic, along with representative examples, is presented in the following section.
@@ 4.2.1 Identified Topics
However, there are some reviews in which users were disappointed due to the late delivery of alerts. An example of such a comment is “… I didn’t receive timely warnings. Alerts can be sent between 10 minutes to an hour late after an earthquake occurs!”
@@
2) Map Usage: The second most-discussed topic concerns map usage, comprising 567 reviews. Common complaints under this topic include insufficient map detail, the clustering of alerts during zooming, and various issues related to functionality and usability. For instance, one user commented, “… The UI is incredibly clunky, maps are extremely slow to load, and warning icons disappear when moving across the map…”
@@
However, some users were satisfied with the map feature in the app. They are happy about aspects such as accuracy of the maps, fastness in loading, and ease of reading and usage. This is shown in the review “… The map feature works very well, and zooming in provides great detail. Positioning the map for a particular global view also works very well…”
@@
3) Monetary Transactions and Ads: This topic contains customer reviews regarding subscriptions, payments, and advertising. Some users give positive ratings and comments because the app is free and does not have ads, such as “Brilliant. This is actual free app… It’s the first 5-star app for me… Plus, there are no ads!”
@@
However, many other users complain about the disruptive ads, for example “… with the latest update, you don’t just get banner ads,but you also get forced full-page pop-up ads, which are incredibly obtrusive and annoying.” Other users also expressed dissatisfaction after purchasing the app or upgrading to the Pro version, stating that the additional features did not meet their expectations. For example, one review noted, “I paid for the Pro version, and there’s nothing special about it. The app doesn’t warn you right away; it has like a 10–15 minute wait time before it pops up. Living in Alaska, this app is worthless to me.”
@@
4) Location Functionality: This topic contains user reviews related to location functionality, such as satisfaction regarding the ability to add their preferred location to be monitored. This is shown in review “I have friends and family scattered all over the country. The app allows me to add the places where they live and provides alerts if there are hazards such as wildfire weather, earthquakes, or hurricanes approaching. It’s easy to customize the alerts I want and to add new places. It’s also great that it will monitor wherever I am.”
@@
Several other reviews mention user dissatisfaction due to inability to add locations or inability to find locations. One such review is “The app will not allow me to add a home location or turn on live monitoring, rendering the entire app useless. My location permissions are set to ’always’, and the app can detect my location via GPS. However, saving never works and instead always produces an error message.”
@@
5) Disaster Updates: This topic includes reviews from users who are mainly pleased to stay informed about disasters occurring in various parts of the world. Some examples are “Great to be able to look all over the world and see what is happening!”, “…The information it contains offers a full view of what earth is doing. Checking out this app is like watching the earth breath and grow or gag and stumble…”. and “… Since I have friends all around the world, I like knowing where the earth is quaking. At times, it is shocking to know how many tremors and earthquakes happen everyday, but I would rather know, than not.”
@@
6) App Crashes: This topic contains user reviews related to the app being unusable due to crashes. Almost all reviews are negative. App crashes can occur in different stages of using the app. One instance is when users attempt to open the application “… lately, either it won’t open at all, or it closes after a couple of seconds…”. It can also happen in other stages of using the app such as “… Sometimes it crashes when I am in the settings, and I have to restart it…”
@@
7) Alert Sound. This topic discusses issues in the level of alert sounds. Sometimes the issue is with the loudness of the alert sound “… but the notification sound is loud and annoying. I turned down my notification volume on my phone, which made other apps quieter, not what I wanted. However, within 2 days, the volume for this app was so high-pitched that it hurts your ears, and the volume on other apps was still low.”. Some users had complaints about the absence of sound, causing users to be unaware if a disaster occurs “A tornado alert was issued by this app, but there are no sounds at all! A wind chime would be more useful than this app. An audible alarm that alerts you should be mandatory on all weather apps.”
@@
8) Slow Performance: This topic contains user reviews related to slow app performance experiences. It can involve slow loading features or slowness when using the features. Example reviews include “Worst app. If you want to view the map and look at details, you have to wait 10 minutes. Super slow.” and “… it’s very slow trying to move around the app, especially when trying to make a watch list…”
@@
9) Bug Fixes: This topic contains customer reviews regarding bug fixing, including both request to fix bugs and appreciation about bugs that have been resolved. Example of a request to fix bugs is “I have not received any notifications from this app for months! I have logged out, uninstalled, and reinstalled multiple times, but still nothing… Please fix it!”. Example of an appreciation about a bug fix is “Thank you very much for fixing the app! It’s the best one for looking up earthquakes and has a great interface and filters!”
@@
10) App Configuration: This topic contains user reviews related to app configuration. Users may express satisfaction with the flexible configuration, or they may encounter problems with the application configuration. Satisfied customer review can look like this “… a fantastic tool. It offers many settings for configurations tailored to specific needs.” Problems in configuration can be “It won’t open settings when I click, and I can’t adjust which warnings or emergencies I want alerts for.”
@@
11) User Interface: This topic contains user reviews related to the app’s interface. Many users may express satisfaction with the UI such as “Great clean new look. Looking forward to future updates, more new features, and improved design.”. They also highlight problematic issues in UI aspects, such as color and size. One review particularly from a user with vision issues stated “The color combination of black with dark blue is terrible. I would have given it five stars if I could actually read some of the menu labels at the bottom of the screen. My eyesight is limited to one eye, with about 65 to 70% vision out of that eye. I need good contrast to be able to see clearly.”
@@
12) User Safety: This topic contains reviews from users who are happy because they feel safe and helped by this application. An example reviews are “… puts a lot of safety ideas for different emergencies in one place. I really like the monitoring feature, with family scattered or traveling a lot” and “Last night we had an earthquake 40km away had this app on my phone unopened for many months but still got a notification seconds before we felt it. It allowed us to immediately recognise what was happening and start heading to a safe area. Quite impressed!”
@@
13) Sign Up and Sign In Problems: This topic contains user reviews complaining about not being able to register or problems during the login process, with almost all reviews negative. Examples are “Keeps saying registration failed… Already missed a fire near me due to this new app. Please fix.” and “It won’t even let me log into the app at all. It keeps saying my login credentials are incorrect, but when I use the same login info on the web browser, it logs right in. I uninstalled it and won’t be reinstalling it!!”
@@
14) Network Connectivity: This topic contains user complaints regarding the inability to use the app due to connectivity issues, almost all reviews negative. For example, as shown in “Keeps telling me no internet connection detected. Yet I can do everything else and use all other apps without any problems.” Another example of a problem is a user failing to connect to the app server “Many times I get an error message about not being able to connect to the server…”. Some users requested the offline mode for the app as shown in review “Great app, although it’s pretty limited without power or internet connectivity. It would be better if it could be mirrored locally for offline use.”
@@
15) Watch Zone Functionality: This topic contains the longest reviews compared to other topics, with the median word count being 27 (see Table 6). This topic contains complaints about the ‘watch zone’ not being able to be set such as “Almost useless. You cannot create watch zones. It just says ’unable to create watch zones’ over and over…”. Besides that, this topic also contains complaints about the watch zone not working properly as shown in review “It gives me notifications every time there’s a thunderstorm anywhere in the state, but when there was actually a fire in my watch zone, it didn’t come up.”
@@
16) Battery Drainage: This topic is quite straightforward, covering the issue of the app consuming a lot of battery, with almost all reviews highly negative. “… the app was using a huge amount of battery and data in the background. Thirty-one percent battery drain was attributed to this app, along with 337MB of data in only a few days of background activity.”
@@
17) Reliability: This topic contains customer reviews regarding whether users find the app reliable or not. Example positive perceptions look like e.g. “The app is very reliable and trustworthy. I depend on it as I live in a quake-prone area, and the information that this app provides is very important to me and my family.” A negative perception is reflected in the review “Unreliable for personal safety and potentially life-threatening during the recent 2020 fires for many people. I don’t know if the networks were jammed or if the app was broken, but we had to rely on looking out a window for a true assessment of whether there are fires near me or not.”
@@
18) Language Preferences: This topic contains customer reviews related to language issues. Many users wish for the app to be available in a language they understand or are fluent, as reflected in the review “… The only criticism I see right now is the lack of supported languages. I think an English version should at least be provided. Other languages could be added by volunteers, for example, through open-source contributions or via a "donate a translation" button in the app.” Some users suggested alternatives for in app translation : “Could you make it possible to have the English translation? Or at least make it possible to copy the text so that I can translate it.”
@@ 4.3 RQ3: Challenges in Disaster Apps
Negative   2663 [24.30%]
Neutral   1911 [17.44%]
Positive   3732 [44.93%]
@@
Negative   2663 [24.30%]
Neutral   1233 [11.25%]
Positive   4410 [53.09%]
Distribution of Review Sentiment: Table 7 presents the distribution of sentiment analysis results across the user review dataset. The proportion of positive sentiment derived from user ratings is 44.93%, while the proportion based on sentiment scores is slightly higher at 53.09%. The percentage of neutral sentiment in user ratings is 17.44%, which is marginally higher than that from sentiment scores at 11.25%. Both methods indicate an identical proportion of negative sentiment at 24.30%.
Table 8: Average rating, sentiment score, and combined sentiment–rating score for each topic.
ID Topic Avg. Sentiment Score Avg. User Rating Avg. Sentiment- User Rating
1 Alert Functionality 3.47 3.48 3.48
2 Map Usage 3.64 3.48 3.56
3 Monetary Transactions and Ads 3.40 3.01 3.21
4 Location Functionality 3.34 2.86 3.10
5 Keep Updates 3.86 4.66 4.26
6 App Crashes 3.24 2.12 2.68
7 Alert Sound 3.33 3.47 3.40
8 Slow Performance 3.27 2.86 3.07
9 Bug Fixes 3.59 2.85 3.22
10 App Configuration 3.18 2.89 3.04
11 User Interface 3.75 3.81 3.78
12 User Safety 3.55 3.71 3.63
13 Sign Up and Sign In Problems 2.94 1.65 2.30
14 Network Connectivity 2.99 2.28 2.64
15 Watch Zone Functionality 3.23 2.53 2.88
16 Battery Drainage 3.25 2.47 2.86
17 Reliability 3.33 3.17 3.25
18 Language Preferences 3.37 2.89 3.13
Topics with Lowest Rating and Sentiment Score: Table 8 presents the average sentiment scores and average user ratings for each topic. The user rating ranges from 1 to 5, while the sentiment score originally ranges from -1 to 1. To make them comparable, the sentiment scores were rescaled to a 1–5 scale using the formula $(x+1)\times 2$. We also added a column called Avg. Sentiment–User Rating, which represents the mean of the sentiment score and user rating. This helps identify topics that users perceive most negatively based on both sentiment and rating. The topic with the lowest combined score is Sign Up and Sign In Problems, with an average score of 2.30. The second lowest is Network Connectivity with 2.64, followed by App Crashes with 2.68. Two other topics fall below the neutral threshold of 3.00, namely Battery Drainage (2.86) and Watch Zone Functionality (2.88). These results suggest that developers aiming to improve disaster applications should prioritize addressing these issues to enhance user satisfaction.
5 Discussions and Recommendations